# Has an AI model been trained on your images?

### v1

@@ Abstract
Keywords Generative AI |||| Membership Inference |||| AI Model Auditing
@@ 1 Introduction
Some advocate for a liberal interpretation of fair-use rules [26] that would allow an AI-powered system to ingest and learn from any web-published content [12]. Others have cried foul, advocating for fair compensation to creators when their content is used to train a generative-AI model, and for the ability to opt out of any training or to have their content “unlearned” by a trained model [18, 27, 30].
As the courts – and the court of public opinion – will no doubt adjudicate these matters in the coming years, the question of whether a model was trained on a specific dataset will be critical. This so-called question of membership inference is challenging for a number of reasons. First, training datasets are massive and often created through large-scale web scraping without careful record keeping [28]. Second, once trained, the models are opaque, making a post-hoc inference challenging. And third, given the competitive landscape and lack of clear laws, there is currently little incentive for rule following, with even some former tech CEOs encouraging young entrepreneurs to steal intellectual property and later hire lawyers to “clean up the mess” [9].
The study of membership inference emerged as deep-learning applications began to be trained on large datasets [11]. Early work focused on membership inference targeting classification (as compared to generative) models [24]. More recently, attention has turned to membership inference for generative models, including large language models (LLMs) [14, 4], generative adversarial networks (GANs) [8, 10], and diffusion-based text-to-image models [29, 3, 31, 13, 32].
@@ 2.1.1 STROLL
This dataset contains $100$ in-training and out-of-training image pairs of outdoor city objects and scenes recorded on a smartphone in the San Francisco Bay area over the course of two days in July 2024. Shown in the top panel of Fig. 1 are representative image pairs (first two rows).
@@ 2.1.2 Carlini
This dataset contains $74$ images that appear to have been memorized [3] by Stable Diffusion (v1.4) [21]. Shown in the middle panel of Fig. 1 are representative images. These images are a tiny fraction of the LAION-5B [23] dataset used to train Stable Diffusion (v1.4). Each in-training image in this dataset is accompanied by its original caption from LAION-5B, which was also used to generate matching out-of-training images using DALL-E (v2) [19] (middle panel of Fig. 1).
@@ 2.1.3 Midjourney
This dataset contains $10$ images that appear to have been memorized [25] by Midjourney (v6) [2]. Shown in the bottom panel of Fig. 1 are representative images. Each in-training image in this dataset is accompanied by its original caption from LAION-5B, which was used to generate matching out-of-training images using Stable Diffusion (v3) [6] (Fig. 1). Unlike the previous two datasets in which the control images were generated using DALL-E, here we use Stable Diffusion (v3) because DALL-E would not generate many of the images in this dataset consisting of recognizable celebrities.
@@ 2.2 Models
We created a custom derivative of the Stable Diffusion (v2.1) variant image model by fine-tuning the 2.1 model weights [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1 on the in-training portion of the STROLL dataset. The official training script [^2] https://github.com/huggingface/diffusers/blob/main/examples/text_to_image/train_text_to_image.py was used to fine-tune the model’s UNet module while keeping the CLIP encoder and variational autoencoder (VAE) weights frozen. The learning rate was set to $10^{-5}$, and the maximum-steps parameter was set to 100100100100333For reference, the fine-tuning of Stable Diffusion v2.1 from v2.0 required $210,000$ steps., while the remaining parameters were left at their recommended default values: image resolution of $512\times 512$, mixed precision turned off, and a random horizontal flip augmentation.
Given a seed image and text prompt as input, the image-to-image feature of this fine-tuned model, with default parameters and varying strengths, was used to power our membership inference. The strength $s_{i}\in[0,1]$ controls the influence of the text prompt relative to the seed image, where a value of 00 yields a generated image that is identical to the seed image, and a value of $1$ generates an image guided fully by the text prompt, effectively ignoring the seed image.
In order to determine the impact of the number of training steps, we created a second fine-tuned model in which the training steps was increased from $100$ to $1{,}000$.
@@
We used the off-the-shelf Stable Diffusion (v1.4) model [^4] https://huggingface.co/CompVis/stable-diffusion-v1-4 and invoked its image-to-image pipeline. All image generation parameters were set to the default values with image strength $s_{i}\in[0,1]$.
@@ 2.2.3 Midjourney/Midjourney (v6)
We used the commercial Midjourney model [^5] https://docs.midjourney.com/docs/image-prompts and manually invoked its image-to-image pipeline through their Discord interface. All parameters were set to the default values, except for the image strength (termed weight in Midjourney), ranging from 00 (yielding a generated image that ignores the seed image) to $3$ (yielding a generated image identical to the seed image). Note that this strength parameter is reversed as compared to Stable Diffusion.
@@ 2.3 Membership Inference
In the second step, the distance $d_{i,j}$ between the seed image $I$ and each generated image $\hat{I}_{i,j}$ is calculated using DreamSim [7]. This perceptual metric of image similarity computes a distance $d_{i,j}\in[0,1]$ where a value of 00 is maximally similar and a value of $1$ is maximally different. For each strength $s_{i}$, the minimum distance across $n$ generated images is retained, yielding a $m$-D vector of distances for each strength value: $\vec{d}=\begin{pmatrix}d_{1}&d_{2}&\ldots&d_{m}\end{pmatrix}$.
@@ 3.1 STROLL
Shown in Fig. 2(a) is a Gaussian fitted log-probability density function to the DreamSim distance between the in-training seed images and the result of image-to-image generation under Stable Diffusion 2.1. Each curve corresponds to a different image-to-image strength parameter $s_{i}$ (see Section 2) where, as strength increases, the seed image has increasingly less impact on the generated image. As expected, for strength parameters close to 00 (thinnest curve), the DreamSim distance is relatively small, and as the strength increases (thicker curves), the distance increases proportionally, meaning that the generated images are increasingly more distinct from the seed image.
@@
Shown in Fig. 2(d) are the same density functions but for the out-of-training seed images. Here we see the same trend, where small strength parameters lead to more similarity as compared to larger strength parameters. However, the mean of these densities as a function of strength $s_{i}$ is larger for these out-of-training images. In particular, notice that the mean of the densities for strengths greater than $0.6$ are significantly larger for out-of-training as compared to in-training images. That is, the images generated with an out-of-training seed are more distinct than those generated with an in-training seed.
@@
Shown in the top panel of Fig. 3 is an in-training seed image (far left) and the resulting image-to-image generation for strengths $s_{i}\in[0.02,0.2,0.4,0.6,0.8,1.0]$. Consistent with the DreamSim distance (Fig. 2(a)), all of the generated images are perceptually similar to the seed image. By comparison, also shown in Fig. 3, is an out-of-training seed image and the resulting image-to-image generation for varying strengths. Again, consistent with the DreamSim distance (Fig. 2(d)), the generated images deviate from the seed starting at a strength of $0.6$.
3.1.1 Alternate Caption
In the above analysis, we assume that the image-to-image generation is provided with the same image caption used in training of the image-generation model. To test the sensitivity of this assumption, alternate captions were generated for each image (see Section 2). Shown in Fig. 2(b) are the same Gaussian fitted log-probability density functions to the DreamSim distances. As compared to the in-training images with the original caption (Fig. 2(a)), the generated images with alternate captions are less similar to their seed images, but still distinct from the out-of-training images (Fig. 2(d)).
3.1.2 Effect Size
3.1.3 Classifier
The distributions in the top panel Fig. 2 show a population-level difference between in-training and out-of-training images. To predict if an individual image belongs to the in-training or out-of-training set, we trained a logistic regression on the $6$-D distance vectors $\vec{d}$ corresponding to the $100$ in-training (original caption), $100$ in-training (alternate caption), $100$ out-of-training and $100$ out-of-training (DALL-E) images. These DALL-E images, generated to match the content of the out-of-training images (see Section 2), balanced the data set for model training. As shown in Fig. 2(e), the distributions for these images are similar to the out-of-training images in panel (d).
3.1.4 Memorization
Shown in Fig. 2(c) is the Gaussian-fitted log probability densities for this model in which we can see a qualitatively different pattern than before, Fig. 2(a). Here, almost regardless of strength, the generated images are uniformly similar to the seed image. This, we posit, is because the prolonged learning caused the model to effectively memorize the training images and associated captions. This is consistent with the results described next.
@@ 3.2 Carlini
As described in Section 2, the Carlini dataset consists of $74$ images used to train Stable Diffusion (v1.4) and, as shown in [3], this model can be coaxed to produce images that are nearly indistinguishable from these training images, Fig. 1.
Shown in Fig. 2(f)-(g) are the DreamSim distances for the $74$ in-training images and $74$ semantically matched out-of-training images (see Section 2). Here, we see the same pattern as with the STROLL results described above: the generated images are perceptually more similar to the in-training seed images than the out-of-training seed images.
Shown in the middle panel of Fig. 3 is an in-training seed image (far left) and the resulting image-to-image generation for varying strengths. As with the previous STROLL results, the generated images are perceptually similar to the seed image. By comparison, the out-of-training seed image yields generated images that deviate from the seed starting at a strength of $0.6$. Interestingly, the out-of-training image at strength $s_{i}=1.0$ is nearly identical to the in-training seed image. This is because, as shown by Carlini et al. [3], this image was memorized during the original training of the model, and so the original prompt yields the training image at a strength of $1.0$ where the seed image is ignored.
@@ 3.2.2 Memorization
Note that the in-training distributions, Fig. 2(f), are qualitatively similar to the distributions in Fig. 2(c) corresponding to the in-training (memorized) results. This, we believe, is because both of these models have memorized some training images and so we see less variation than in the case when the model was simply exposed to these images.
@@ 3.3 Midjourney
As described in Section 2, the Midjourney dataset consists of $10$ images that appear to have been part of the training datatset for Midjourney (v6). In particular, as shown in [25], Midjourney can be coaxed to produce images that are nearly indistinguishable from these well-recognized images, Fig. 1.
Shown in Fig. 2(h)-(i) are the DreamSim distances for the $10$ in-training images and $10$ semantically matched out-of-training images (see Section 2). Here, we see the same pattern as with the STROLL and Carlini results: the generated images are perceptually more similar to the in-training seed images than the out-of-training seed images.
Shown in the bottom panel of Fig. 3 is an in-training seed image (far left) and the resulting image-to-image generation for varying strengths. As with the previous STROLL and Carlini results, the generated images are perceptually similar to the seed image. By comparison, the out-of-training seed image yields generated images that deviate from the seed.
We again see that the out-of-training image, at a strength $s_{i}=0$ where the seed image is ignored, is nearly identical to the in-training seed image. This is because this image was memorized during the original training of the model [25], and so the original prompt yields the training image.
@@ 3.3.1 Memorization
As before, the in-training distributions, Fig. 2(h), are qualitatively similar to the distributions in Fig. 2(c) and (f) corresponding to the STROLL in-training (memorized) and Carlini in-training results. This, again, is because all three of these models have memorized some training images.

### v2

# GenAI Confessions: Black-box Membership Inference for Generative Image Models

@@ Abstract
Figure 1: Overview of Our Membership Inference Method. Images are generated using an image-to-image pipeline, seeded with an in-training image (bottom) and an out-of-training image (top), across increasing strengths. A larger strength corresponds to reduced emphasis on the seed image. The plot on the left shows the similarity between each generated image and its corresponding seed image (the DreamSim image similarity metric), where smaller values indicate greater similarity. The in-training seed consistently yields more similar outputs than the out-of-training seed. Our membership inference method exploits this phenomenon.
Project Page:          genai-confessions.github.io
Data:                    hf.co/datasets/faridlab/stroll
@@ 1 Introduction
As the courts – and the court of public opinion – adjudicate these matters in the coming years, the question of whether a model was trained on a specific dataset will be critical. This so-called question of membership inference is challenging for a number of reasons. First, training datasets are massive and often created through large-scale web scraping without careful record keeping [25]. Second, once trained, the models are opaque, making a post-hoc inference challenging. And third, given the competitive landscape and lack of clear laws, there is currently little incentive for rule following, with even some former tech CEOs encouraging young entrepreneurs to steal intellectual property and later hire lawyers to “clean up the mess” [10].
The study of membership inference emerged as deep-learning applications started to be trained on large datasets [12]. Early work focused on membership inference targeting classification (as compared to generative) models [23]. More recently, attention has turned to membership inference for generative models, including large language models (LLMs) [14, 4], generative adversarial networks (GANs) [9, 11], and diffusion-based text-to-image models [26, 3, 27, 13, 28].
@@
The contributions of this paper can be summarized as follows:
  • A computationally efficient and easy to implement, black-box membership inference method for generative-AI image models (Figure 1);
  • A dataset (STROLL) of semantically matched image pairs for evaluating membership inference;
  • Empirical analysis of membership inference and memorization across different model architectures.
@@ 2.1.1 STROLL
This dataset contains $100$ in-training and out-of-training image pairs of outdoor city objects and scenes recorded on a smartphone in the San Francisco Bay area over the course of two days in July 2024. Shown in the top panel of Figure 2 are representative image pairs (first two rows).
@@ 2.1.2 Carlini
This dataset contains $74$ images that appear to have been memorized [3] by Stable Diffusion (v1.4) [20]. Shown in the middle panel of Figure 2 are representative images. These images are a tiny fraction of the LAION-5B [22] dataset used to train Stable Diffusion (v1.4). Each in-training image in this dataset is accompanied by its original caption from LAION-5B, which was also used to generate matching out-of-training images using DALL-E (v2) [18] (middle panel of Figure 2).
@@ 2.1.3 Midjourney
This dataset contains $10$ images that appear to have been memorized [24] by Midjourney (v6) [2]. Shown in the bottom panel of Figure 2 are representative images. Each in-training image in this dataset is accompanied by its original caption from LAION-5B, which was used to generate matching out-of-training images using Stable Diffusion (v3) [7] (Figure 2). Unlike the previous two datasets in which the control images were generated using DALL-E, here we use Stable Diffusion (v3) because DALL-E would not generate many of the images in this dataset consisting of recognizable celebrities.
@@ 2.2 Models
We created a custom derivative of the Stable Diffusion (v2.1) variant image model by fine-tuning the 2.1 model weights111www.huggingface.co/stabilityai/stable-diffusion-2-1 on the in-training portion of the STROLL dataset. The official training script [^2] https://github.com/huggingface/diffusers/blob/main/examples/text_to_image/train_text_to_image.py was used to fine-tune the model’s UNet module while keeping the CLIP encoder and variational autoencoder (VAE) weights frozen. The learning rate was set to $10^{-5}$, and the maximum-steps parameter was set to 100100333For reference, fine-tuning SD v2.1 from v2.0 took 210,000210{,}000 steps., while the remaining parameters were left at their recommended default values: image resolution of $512\times 512$, mixed precision turned off, and a random horizontal flip augmentation.
Given a seed image and text prompt as input, the image-to-image feature of this fine-tuned model, with default parameters and varying strengths, was used to power our membership inference. The strength $s_{i}\in[0,1]$ controls the influence of the text prompt relative to the seed image, where a value of 0 yields a generated image that is identical to the seed image, and a value of $1$ generates an image guided fully by the text prompt, effectively ignoring the seed image.
In order to determine the impact of the number of training steps, we created a second fine-tuned model in which the training steps was increased from $100$ to $1{\tiny,}000$.
@@
We used the off-the-shelf Stable Diffusion (v1.4) model444www.huggingface.co/CompVis/stable-diffusion-v1-4 and invoked its image-to-image pipeline. All image generation parameters were set to the default values with image strength $s_{i}\in[0,1]$.
@@ 2.2.3 Midjourney/Midjourney (v6)
We used the commercial Midjourney model555docs.midjourney.com/docs/image-prompts and manually invoked its image-to-image pipeline through their Discord interface. All parameters were set to the default values, except for the image strength (termed weight in Midjourney), ranging from 0 (yielding a generated image that ignores the seed image) to $3$ (yielding a generated image identical to the seed image). Note that this strength parameter is reversed as compared to Stable Diffusion.
@@ 2.3 Membership Inference
In the second step, the distance $d_{i,j}$ between the seed image $I$ and each generated image $\hat{I}_{i,j}$ is calculated using DreamSim [8]. This perceptual metric of image similarity computes a distance $d_{i,j}\in[0,1]$ where a value of 0 is maximally similar and a value of $1$ is maximally different. For each strength $s_{i}$, the minimum distance across $n$ generated images is retained, yielding a $m$-D vector of distances for each strength value: $\vec{d}=\begin{pmatrix}d_{1}&d_{2}&\ldots&d_{m}\end{pmatrix}$.
@@ 3.1 STROLL
Shown in Figure 3(a) is a Gaussian fitted log-probability density function to the DreamSim distance between the in-training seed images and the result of image-to-image generation under Stable Diffusion 2.1. Each curve corresponds to a different image-to-image strength parameter $s_{i}$ (see Section 2) where, as strength increases, the seed image has increasingly less impact on the generated image. As expected, for strength parameters close to 0 (thinnest curve), the DreamSim distance is relatively small, and as the strength increases (thicker curves), the distance increases proportionally, meaning that the generated images are increasingly more distinct from the seed image.
@@
Shown in Figure 3(d) are the same density functions but for the out-of-training seed images. Here we see the same trend, where small strength parameters lead to more similarity as compared to larger strength parameters. However, the mean of these densities as a function of strength $s_{i}$ is larger for these out-of-training images. In particular, notice that the mean of the densities for strengths greater than $0.6$ are significantly larger for out-of-training as compared to in-training images. That is, the images generated with an out-of-training seed are more distinct than those generated with an in-training seed.
@@
Shown in the top panel of Figure 4 is an in-training seed image (far left) and the resulting image-to-image generation for strengths $s_{i}\in[0.02,0.2,0.4,0.6,0.8,1.0]$. Consistent with the DreamSim distance (Figure 3(a)), all of the generated images are perceptually similar to the seed image. By comparison, also shown in Figure 4, is an out-of-training seed image and the resulting image-to-image generation for varying strengths. Again, consistent with the DreamSim distance (Figure 3(d)), the generated images deviate from the seed starting at a strength of $0.6$.
Independent-samples, two-sided t-tests reveal a significant difference between the DreamSim distributions for the in-training vs out-of-training data at a strength $s_{i}\geq 0.4$, with an increasing effect size (Cohen’s D) with increasing strength:
  • $0.02$ (t(198)=0.5,p=0.6,D=0.07t(198)=0.5,~p=0.6,~D=0.07)
  • $0.2$ (t(198)=2.3,p=0.02,D=0.32t(198)=2.3,~p=0.02,~D=0.32)
  • $0.4$ ($t(198)=7.3,~p<10^{-8},~D=1.0$)
  • $0.6$ ($t(198)=11.2,~p<10^{-8},~D=1.6$)
  • $0.8$ ($t(198)=14.5,~p<10^{-8},~D=2.1$)
  • $1.0$ ($t(198)=14.8,~p<10^{-8},~D=2.1$)
@@
3.1.1 Alternate Caption
In the above analysis, we assume that the image-to-image generation is provided with the same image caption used in training of the image-generation model. To test the sensitivity of this assumption, alternate captions were generated for each image (see Section 2). Shown in Figure 3(b) are the same Gaussian fitted log-probability density functions to the DreamSim distances. As compared to the in-training images with the original caption (Figure 3(a)), the generated images with alternate captions are less similar to their seed images, but still distinct from the out-of-training images (Figure 3(d)).
An independent-samples, two-sided t-tests again reveals a significant difference between the DreamSim distributions for the in-training vs out-of-training data at a strength $s_{i}\geq 0.4$:
  • $0.02$ (t(198)=0.7,p=0.5,D=0.1t(198)=0.7,~p=0.5,D=0.1)
  • $0.2$ (t(198)=1.6,p=0.1,D=0.2t(198)=1.6,~p=0.1,D=0.2)
  • $0.4$ ($t(198)=5.7,~p<10^{-8},~D=0.8$)
  • $0.6$ ($t(198)=8.5,~p<10^{-8},~D=1.2$)
  • $0.8$ ($t(198)=11.1,~p<10^{-8},~D=1.6$)
  • $1.0$ ($t(198)=10.5,~p<10^{-8},~D=1.5$)
3.1.2 Effect Size
3.1.3 Classifier
The distributions in the top panel Figure 3 show a population-level difference between in-training and out-of-training images. To predict if an individual image belongs to the in-training or out-of-training set, we trained a logistic regression on the $6$-D distance vectors $\vec{d}$ corresponding to the $100$ in-training (original caption), $100$ in-training (alternate caption), $100$ out-of-training and $100$ out-of-training (DALL-E) images. These DALL-E images, generated to match the content of the out-of-training images (see Section 2), balanced the data set for model training. As shown in Figure 3(e), the distributions for these images are similar to the out-of-training images in panel (d).
3.1.4 Memorization
Shown in Figure 3(c) is the Gaussian-fitted log probability densities for this model in which we can see a qualitatively different pattern than before, Figure 3(a). Here, almost regardless of strength, the generated images are uniformly similar to the seed image. This, we posit, is because the prolonged learning caused the model to effectively memorize the training images and associated captions. This is consistent with the results described next.
@@ 3.2 Carlini
As described in Section 2, the Carlini dataset consists of $74$ images used to train Stable Diffusion (v1.4) and, as shown in [3], this model can be coaxed to produce images that are nearly indistinguishable from these training images, Figure 2.
Shown in Figure 3(f)-(g) are the DreamSim distances for the $74$ in-training images and $74$ semantically matched out-of-training images (see Section 2). Here, we see the same pattern as with the STROLL results described above: the generated images are perceptually more similar to the in-training seed images than the out-of-training seed images.
Shown in the middle panel of Figure 4 is an in-training seed image (far left) and the resulting image-to-image generation for varying strengths. As with the previous STROLL results, the generated images are perceptually similar to the seed image. By comparison, the out-of-training seed image yields generated images that deviate from the seed starting at a strength of $0.6$. Interestingly, the out-of-training image at strength $s_{i}=1.0$ is nearly identical to the in-training seed image. This is because, as shown by Carlini et al. [3], this image was memorized during the original training of the model, and so the original prompt yields the training image at a strength of $1.0$ where the seed image is ignored.
@@ 3.2.2 Memorization
Note that the in-training distributions, Figure 3(f), are qualitatively similar to the distributions in Figure 3(c) corresponding to the in-training (memorized) results. This, we believe, is because both of these models have memorized some training images and so we see less variation than in the case when the model was simply exposed to these images.
@@ 3.3 Midjourney
As described in Section 2, the Midjourney dataset consists of $10$ images that appear to have been part of the training datatset for Midjourney (v6). In particular, as shown in [24], Midjourney can be coaxed to produce images that are nearly indistinguishable from these well-recognized images, Figure 2.
Shown in Figure 3(h)-(i) are the DreamSim distances for the $10$ in-training images and $10$ semantically matched out-of-training images (see Section 2). Here, we see the same pattern as with the STROLL and Carlini results: the generated images are perceptually more similar to the in-training seed images than the out-of-training seed images.
Shown in the bottom panel of Figure 4 is an in-training seed image (far left) and the resulting image-to-image generation for varying strengths. As with the previous STROLL and Carlini results, the generated images are perceptually similar to the seed image. By comparison, the out-of-training seed image leads to images that deviate more noticeably.
We again see that the out-of-training image, at a strength $s_{i}=0$ where the seed image is ignored, is nearly identical to the in-training seed image. This is because this image was memorized during the original training of the model [24], and so the prompt simply reproduces it.
@@ 3.3.1 Memorization
As before, the in-training distributions, Figure 3(h), are qualitatively similar to those in Figure 3(c) and (f) corresponding to the STROLL in-training (memorized) and Carlini in-training results. This, again, is because all three of these models have memorized some training images.
@@ 3.3.2 Classifier
3.4 Comparison to Previous Work
Comparing membership inference methods for generative-AI image models remains challenging due to the lack of standardized benchmarks and varying problem definitions/configurations (including the frequency and intensity with which images are presented to the model during training). Nonetheless, we compare our method to existing membership inference approaches for the latest generative-AI image model architectures (diffusion) in terms of computational demands and overall effectiveness.
The method in [26] assumes access to a confirmed subset of the targeted model’s in-training and out-of-training data. Its reported accuracy ranges from $65\%$ to $100\%$ across multiple models and datasets. The method in [27] assumes access to internal representations of the model during diffusion steps. At a fixed false positive rate of $1\%$, it achieves a true positive rate ranging from $54\%$ to $68\%$ across multiple datasets. Similarly, the method in [13] also relies on internal representations during diffusion steps. At a fixed false positive rate of $1\%$, it achieves a true positive rate between $50\%$ and $58\%$ across different datasets. And, the method in [5] is an ensemble of four other methods, whose predictions are analyzed for statistical significance and overall confidence. At a fixed false positive rate of $1\%$, it achieves a true positive rate ranging from $25\%$ to $100\%$ across multiple datasets.
By contrast, our method does not require access to a confirmed subset of in-training/out-of-training data or internal representations from diffusion steps and achieves an average accuracy between $85\%$ and $93\%$. At a fixed false positive rate of $1\%$, our method attains an average true positive rate of between $74\%$ and $90\%$.
When tested in the wild, many existing methods have been found largely ineffective [4]. We, on the other hand, demonstrate that the DreamSim trends leveraged by our method persist even in an in-the-wild setting with Midjourney (v6) (see Section 3.3).
Most effective membership inference and training data extraction methods are computationally expensive. The method in [3], which assumes a white-box scenario with additional access to the model’s internals, achieves a true positive rate of $71\%$ at a false positive rate of $1\%$. This approach, however, is computationally demanding, requiring training $16$ shadow models, each of which computes loss for all known in-training data points at each of $1{\small,}000$ diffusion steps. By contrast, our method does not require training and only needs to perform six generations with at most $50$ diffusion steps.